An unmanned research submarine desecends to 150 feet below the surface of the ocean. Then it continues to go deeper, taking a water sample every 25 feet. What is the depth when it takes the 15th sample?
Answer by stanbon(75887)

An unmanned research submarine desecends to 150 feet below the surface of the ocean. Then it continues to go deeper, taking a water sample every 25 feet. What is the depth when it takes the 15th sample?
------
Equation:
depth = 150 + 25x
d(15) = 150 + 25*15 = 525 ft
===============================
Cheers,
Stan H.